OpenStudy (anonymous):

find the critical numbers of f determine the intervals on which the function is increasing or decreasing, and identify all relative extrema. f(x)=(x-1)^2 (x+2)

OpenStudy (anonymous):

@factor can you help me with this?

OpenStudy (anonymous):

f'(x) = 2(x-1)(x+2)+(x-1)^2

OpenStudy (anonymous):

so critical numbersa are x=1 and x=-2 right?

OpenStudy (ranga):

f'(x) = 2(x-1)(x+2)+(x-1)^2 = (x-1)(2x+4+x-1) = (x-1)(3x+3) f'(x) = 3(x-1)(x+1) The critical points are: x = 1 and x = -1

OpenStudy (ranga):

Analyze f'(x) in the intervals: (-infinity, -1), (-1, 1) and (1, infinity) f'(x) = 3(x-1)(x+1) f'(-2) = (-ve)(-ve) = +ve. Therefore function is increasing in the interval (-infinity, -1). f'(0) = (-ve)(+ve) = -ve Therefore function is decreasing in the interval (-1, 1). f'(2) = (+ve)(+ve) = +ve Therefore function is increasing in the interval (1, infinity). f(x) attains a maximum at x = -1 and a minimum at x = +1.

OpenStudy (anonymous):

Thank you

OpenStudy (ranga):

You are welcome.